# Efficient & Correct Predictive Equivalence for Decision Trees

### v2

@@ 2.4 Logic-Based Explanations
Recent years witnessed the emergence of rigorous, logic-based, explanability. This subsection provides a brief account of the progress in this field of research.
@@
For a given classification model $M$, we define the set of AXps for an instance $(\mathbf{v},c)$ by $\mathbb{A}((\mathbf{v},c);M)=\{{\mathcal{A}}\subseteq\mathbb{L}\,|\,\mathsf{AXp}({\mathcal{A}};(\mathbf{v},c),M)\}$. Furthermore, we observe that, given the definitions above, the predicate $\mathsf{WAXp}$ does not depend on $\mathbf{v}$, only on some partial assignment (that may be extracted from $\mathbf{v}$). As a result, when convenient, we will use the representation $\mathsf{WAXp}({\mathcal{A}};c,M)$ that holds true for partial assignments such that (4) holds, when $M$ computes classification function $\kappa$ and predicts $c\in\mathbb{K}$. This alternative representation will also be used to represent AXps. Thus, we can now define the set of all partial assignments that represent (W)AXps for a given prediction $c$:
@@
Furthermore, each AXp is uniquely defined by a partial assignment and a prediction. We can then talk about the sets of pairs, of partial assignments and predictions, such that the partial assignments are either sufficient for some prediction $c$,
@@ 2.5 Additional Concepts
Let $\mathbb{S}(P,I)$ represent the set of solutions for some computational problem $P$ on input $I$. For example, in the case of the QM method, if $P$ represents the QM method, and $I$ denotes some boolean function, then any minimum-cost DNF would be included in $\mathbb{S}(P)$. An algorithm $A$ for problem $P$ on input $I$ implements some form of non-determinism if multiple executions of $A$ can produce different elements of $\mathbb{S}(P)$. If any execution of algorithm $A$ for problem $P$ on input $I$ always computes the same solution $S\in\mathbb{S}(P,I)$, then $A$ is deterministic. Non-determinism may represent an one or more of an algorithm’s design decisions, e.g. the use of randomization, or it may manifest itself by the specificities of the programming language used for implementing the algorithm, e.g. the implementation of sorting not being stable, or the use of hashes returning elements in some arbitrary order. Most algorithms for solving combinatorial problems implement some sort of non-determinism (Gomes et al., 1998, 2000; Biere et al., 2021). In addition, the problem of non-determinism in sequential programs is the subject of continued research (Mudduluru et al., 2021; Miao et al., 2025).
3 Assessment of the MBDSR Approach
3.1 MBDSR’s Worst-Case Exponential Time & Space
@@ 3.2 Potential Unsoundness of QM-based Predictive Equivalence
As stated in Section˜1, the MBDSR algorithm for predictive equivalence proposed in (McTavish et al., 2025a, c) can yield incorrect results if off-the-shelve implementations of the QM method are used. This observation also explains why McTavish et al. impose a requirement of deterministic set covering step of the QM method.) In general, if two DTs are not predictive equivalent, then any implementation of the QM method will correctly report non-equivalence. However, if two DTs are predictive equivalent, then the QM method is not in general guaranteed to report equivalence. The reason for why the use of the QM method for predictive equivalence (McTavish et al., 2025c, a) can be problematic is rather simple.
The Blake canonical form is (as the name indicates) canonical (Blake, 1937; Brown, 1990). Canonicity signifies that there exists a single unique representation for some boolean function (Bryant, 1986; Wegener, 2000; Crama and Hammer, 2011). However, if a representation is not canonical, then there may exist multiple distinct representations for a given boolean function. A potential problem of McTavish et al.’s algorithm is that it computes a minimum-cost (e.g. number of literals or terms) representation of some boolean function, and this is not a canonical representation. [^8] Evidently, the concept of canonicity is fundamental for ensuring uniqueness of representations (Bryant, 1986; Wegener, 2000; Crama and Hammer, 2011). The previous work of McTavish et al. (McTavish et al., 2025c, a) does not discuss the possible issues resulting from adopting the QM method. However, the description of the proposed predictive equivalence algorithm overcomes the limitation of the QM method by imposing that the set covering step of the QM method be deterministic. [^9] Although the works of McTavish et al. (McTavish et al., 2025c, a) do not clarify the implementation of the QM method used, analysis of the code confirms it is SymPy (Meurer et al., 2017). Furthermore, the authors report analyzing the code of SymPy and testing its results, aiming at ensuring that the execution is deterministic (McTavish et al., 2025d). Nevertheless, an elegant and simpler solution to overcome the difficulties of the QM method, would be to compare DTs based on the BCF, since this is a canonical representation (Blake, 1937; Brown, 1990). This would not require imposing any constraint on the QM method used.
This section provides additional detail demonstrating the above arguments, namely the issues with the unrestricted used of the QM method. Example˜3 analyzes an a boolean function with two distinct minimum cost DNFs. If implementations of QM use any form of non-determinism, i.e. they may not not always report the same minimum-cost DNF, then the use of QM for predictive equivalence, as proposed in the MBDSR approach, can declare predictive equivalent DTs not to be equivalent.
@@
Furthermore, again by inspection of the Karnaugh maps, we can conclude that both $\mathbb{I}_{1}$ and $\mathbb{I}_{2}$ are minimum-cost DNF representations of $f$, as the two maps confirm. The set of prime implicants $\mathbb{I}_{1}$ covers all the minterms (in red), and the set of prime implicants $\mathbb{I}_{2}$ also covers all the minterms (in blue), i.e. both are minimum-cost DNFs (for some class). Either $\mathbb{I}_{1}$ or $\mathbb{I}_{2}$ could be reported by an implementation of the QM method. This holds true as long as the implementation of the QM method implements any form of non-determinism.
The previews arguments justify the following result.
Theorem 2
Regarding the MBDSR algorithm for deciding predictive equivalence:
@@
    MBDSR is incorrect for implementations of the QM method that exhibit non-determinism.
@@
    MBDSR is correct for implementations of the QM method that do not exhibit non-determinism.
As stated earlier, algorithms for solving combinatorial problems often implement some form of non-determinism, either deliberate as part of the algorithm itself or motivated by the programming language used (e.g. recent references on the topic of non-determinism in sequential programs include (Mudduluru et al., 2021; Miao et al., 2025)). Furthermore, it should also be underscored that the presentation in McTavish et al.’s work (McTavish et al., 2025c, a) does not address any of the issues resulting from the QM method producing a representation that is not canonical, apart from requiring that a minimal set cover to be computed deterministically.
@@ 4 Polynomial-Time Algorithms
This section shows that the computational overhead of the MBDSR approach can be avoided. Concretely, this signifies that the computation of $\mathsf{BCF}{{\mathcal{T}}}$ and $\mathcal{T}_{\mathsf{DNF}}$ (and so the execution of the QM method) is unnecessary for the computational problems studied in McTavish et al.’s work (McTavish et al., 2025c, a). To prove this claim, we show that each algorithm proposed in (McTavish et al., 2025c, a) that uses $\mathcal{T}_{\mathsf{DNF}}$ can be replaced by an algorithm with a running time polynomial in the size of the original DT.
@@ 4.1 Completeness & Succinctness
Given the above, a partial assignment $\mathcal{A}$} is a WAXp of DT $\mathcal{T}$} if and only if it is also a WAXp of any DT $\mathcal{T}$$^{\prime}$ that is predictive equivalent to $\mathcal{T}$}. Furthermore, prediction in the presence of missing data is tightly related with a partial assignment representing a WAXp.
@@
As noted in recent work (Izza et al., 2022; Audemard et al., 2022b; McTavish et al., 2025c, a), decision trees can exhibit path literal redundancy, and so the explanation for a DT prediction can be smaller, possibly much smaller, than the path size in a DT. McTavish et al.’s propose to computing such explanations (and so deciding predictions given missing data) using ${\mathcal{T}}_{\mathsf{DNF}}$. However, as demonstrated in Section˜3 this is worst-case exponential in time and space. In contrast, earlier work proposed a polynomial-time algorithm for explaining an instance $(\mathbf{v},c)$ (e.g. (Izza et al., 2020, 2022; Audemard et al., 2022b)). Here, we opt instead for explaining a partial assignment that is known to be a WAXp for some class $c$. Algorithm˜4 summarizes the computation of one AXp, which requires that the starting partial assignment is an WAXp for class $c$.
@@
The algorithm’s invariant is that the working partial assignment $\mathcal{X}$} is a WAXp, from which features are removed which the invariant is preserved. Due to monotonicity of the definition of WAXp, the partial assignment that is returned represents an AXp. Clearly, the algorithm ensures the property of succinctness of explanations, and runs in polynomial time in the size of the DT. Algorithm˜4 mimics the algorithm proposed in earlier work (Izza et al., 2022), but adapted to working with partial assignments instead of a concrete instance.
@@ 5 Experimental Evidence
The work of McTavish et al. (McTavish et al., 2025c, a) offers comprehensive evidence regarding the importance of eliminating predictive equivalent DTs from the Rashomon set. Also, the algorithms proposed in the previous section exhibit negligible running times, even for DTs of fairly large size. (Our results concur with similar results were obtained in earlier work when computing instance-specific explanations (Izza et al., 2022).) As a result, the focus of our experiments is to demonstrate the practical effect of exercising the worst-case running time of the QM method. Since the proof Theorem˜1 only considers a single instance, the experiments target the complete compilation of $\mathsf{BCF}$ for different values of $r$.
@@ Results.
Finally, we underscore that the experimental results consider a special DT construction aiming at highlighting the performance limitations of using an implementation of the QM method. However, DTs of large size can occur in some high-risk (Rudin, 2019) application domains, where high accuracy is paramount. For example, (Ghiasi et al., 2020) reports a DT with 161 nodes and 20 real-valued features. In those cases, the use of the QM method will also not scale, even if high-performance solutions are exploited (Coudert, 1994). (We note that SymPy is known to be inefficient for formulas with more than 8 boolean variables. [^14] This is discussed for example in https://docs.sympy.org/latest/modules/logic.html.)
@@ 6.2 Predictive Equivalence and Corrected Measures of Importance
It is well-known that SHAP scores approximate a theoretical definition of Shapley values for XAI (Strumbelj and Kononenko, 2010, 2014). Thus, the quality of the approximation provided by SHAP scores depends on a number of factors. However, recent work proved that the proposed definition of Shapley values for XAI (Strumbelj and Kononenko, 2010, 2014; Lundberg and Lee, 2017) can mislead human decision makers (Marques-Silva and Huang, 2024; Huang and Marques-Silva, 2024), by assigning no importance to features that are critical for a given prediction, and by assigning some importance to features that bear no influence in that prediction. Since it is unclear when Shapley values for XAI (and so SHAP scores) can mislead, the use of the existing definition of Shapley values for XAI should be discouraged.
@@
As argued earlier in this section, in practice appears harder to compute the Shapley values for every instance than simply deciding predictive equivalence. However, Theorem˜10 shows the connections that logic-based XAI reveals. Nevertheless, Theorem˜10 is one example of several measures of importance that can be considered:
@@ 6.3 Complexity Results Beyond DTs
Regarding predictive equivalence, this paper (see Algorithm˜5) proves that the problem is in P for decision trees. The same algorithm can be adapted for tractable fragments of multivariate decision trees (Carbonnel et al., 2023; Cooper, 2025). While in general the comparison of non-canonical representations for predictive equivalence can be unsound, this is a correct approach for when classifiers exhibit canoninicity. This is well-known to be the case with $\{\mathsf{BCF}_{c}(C)\,|\,c\in{\mathcal{K}}\}$, i.e. the sets of prime implicants for each predicted class for classifier $C$ (Blake, 1937; Brown, 1990; Darwiche and Marquis, 2002). In addition, it is well-known that reduced ordered binary decision diagrams (ROBDDs) (Bryant, 1986; Darwiche and Marquis, 2002) and sentential decision diagrams (SDDs) (Darwiche, 2011) are canonical representations (Bryant, 1986; Darwiche and Marquis, 2002; Van den Broeck and Darwiche, 2015). As a result, predictive equivalence can be decided in polynomial-time for classifiers represented with ROBDDs and SDDs.

### v3

@@ 2.4 Logic-Based Explanations
Recent years witnessed the emergence of rigorous, logic-based, explanability Marques-Silva (2022); Marques-Silva and Ignatiev (2022); Darwiche (2023). This subsection provides a brief account of the progress in this field of research.
@@
For a given classification model $M$, we define the set of AXps for an instance $(\mathbf{v},c)$ by $\mathbb{A}((\mathbf{v},c);M)=\{{\mathcal{A}}\subseteq\mathbb{L}\,|\,\mathsf{Consistent}({\mathcal{A}})\land\mathsf{AXp}({\mathcal{A}};(\mathbf{v},c),M)\}$. Furthermore, we observe that, given the definitions above, the predicate $\mathsf{WAXp}$ does not depend on $\mathbf{v}$, only on some partial assignment (that may be extracted from $\mathbf{v}$). As a result, when convenient, we will use the representation $\mathsf{WAXp}({\mathcal{A}};c,M)$ that holds true for partial assignments such that (4) holds, when $M$ computes classification function $\kappa$ and predicts $c\in\mathbb{K}$. This alternative representation will also be used to represent AXps. Thus, we can now define the set of all partial assignments that represent (W)AXps for a given prediction $c$:
@@
Furthermore, each AXp is uniquely defined by a partial assignment and a prediction. We can then talk about the sets of pairs, of partial assignments and predictions, such that the partial assignments are sufficient for some prediction $c$,
@@ 2.5 Additional Concepts
Let $\mathbb{S}(P,I)$ represent the set of solutions for some computational problem $P$ on input $I$. For example, in the case of the QM method, if $P$ represents the QM method, and $I$ denotes some boolean function, then any minimum-cost DNF would be included in $\mathbb{S}(P)$. An algorithm $A$ for problem $P$ on input $I$ implements some form of non-determinism if multiple executions of $A$ can produce different elements of $\mathbb{S}(P)$. If any execution of algorithm $A$ for problem $P$ on input $I$ always computes the same solution $S\in\mathbb{S}(P,I)$, then $A$ is deterministic. Non-determinism may represent one or more of an algorithm’s design decisions, e.g. the use of randomization, or it may manifest itself by the specificities of the programming language used for implementing the algorithm, e.g. the implementation of sorting not being stable, or the use of hashes returning elements in some arbitrary order. Most algorithms for solving combinatorial problems implement some sort of non-determinism (Gomes et al., 1998, 2000; Biere et al., 2021). In addition, the problem of non-determinism in sequential programs is the subject of continued research (Mudduluru et al., 2021; Miao et al., 2025).
3 Assessment of McTavish et al.’s Approach
3.1 Worst-Case Exponential Time & Space
@@ 3.2 Potential Unsoundness of QM-based Predictive Equivalence
This section details why the adoption of the QM method for predictive equivalence can yield incorrect results. Section˜3.3 briefly discusses the constraints imposed in the work of McTavish et al. to circumvent this problem. [^8] This discussion is included because the previous work of McTavish et al. (McTavish et al., 2025c, a) does not discuss the possible issues resulting from adopting the QM method, neither does it explain the need for the constraints imposed on the QM method used.
In general, if two DTs are not predictive equivalent, then any implementation of the QM method will correctly report non-equivalence. However, if two DTs are predictive equivalent, then the QM method is not in general guaranteed to report equivalence. The reason for why the use of the QM method for predictive equivalence (Quine, 1952, 1955; McCluskey, 1956) can be problematic is rather simple. The Blake canonical form is (as the name indicates) canonical (Blake, 1937; Brown, 1990). Canonicity signifies that there exists a single unique representation for some boolean function (Bryant, 1986; Wegener, 2000; Crama and Hammer, 2011). However, if a representation is not canonical, then there may exist multiple distinct representations for a given boolean function. A potential problem of adopting the QM methods for predictive equivalence is that it computes a minimum-cost (e.g. number of literals or terms) representation of some boolean function, and this is not a canonical representation. [^9] Evidently, the concept of canonicity is fundamental for ensuring uniqueness of representations (Bryant, 1986; Wegener, 2000; Crama and Hammer, 2011).
This section provides additional detail demonstrating the above arguments, namely the issues with the unrestricted used of the QM method. Example˜3 analyzes an a boolean function with two distinct minimum cost DNFs. If implementations of QM use any form of non-determinism, i.e. they may not not always report the same minimum-cost DNF, then the use of QM for predictive equivalence can declare predictive equivalent DTs not to be equivalent.
@@
Furthermore, again by inspection of the Karnaugh maps, we can conclude that both $\mathbb{I}_{1}$ and $\mathbb{I}_{2}$ are minimum-cost DNF representations of $f$, as the two maps confirm. The set of prime implicants $\mathbb{I}_{1}$ covers all the minterms (in red), and the set of prime implicants $\mathbb{I}_{2}$ also covers all the minterms (in blue), i.e. both are minimum-cost DNFs (for some class). Either $\mathbb{I}_{1}$ or $\mathbb{I}_{2}$ could be reported by an implementation of the QM method. This holds true as long as the execution of the QM method is not guaranteed to be deterministic.
3.3 Use of QM in the Approach of McTavish et al.
The description of predictive equivalence in McTavish et al. (2025c, a) imposes two constraints on the execution of QM: (i) the prime implicants must be sorted (and it is implicitly assumed that the truth table with the minterms is also sorted); and (ii) the execution of the second phase of the QM method, i.e. the set-covering step, must be deterministic. McTavish et al. (2025c, a) do not justify the need for these two constraints. Nevertheless, with the explicit assumption that these two constraints hold, then McTavish et al. (2025c, a) prove that predictive equivalence will be decided correctly.
We underscore that predictive equivalence could be decided correctly without the second constraint, because BCF is a canonical representation. Hence, just sorting and comparing the lists of prime implicants would suffice to decide predictive equivalence. The bottom line is that there is not formal need for using the set-covering step of the QM method and, with that, imposing the need for deterministic execution. (Evidently, the set-covering step can be important from the perspective of interpretability.)
Furthermore, ensuring determinism can be a challenge in practice. This depends on whether or not non-deterministic choices are made by the set-covering algorithm, but also on whether the programming language used, and the programming constructs used in the implementation of the set-covering algorithm, guarantee determinism. For some programming languages it is known that is not the case Mudduluru et al. (2021); Miao et al. (2025). (McTavish et al., 2025c, a) do not detail the implementation of the QM method used, nor discuss whether or not it guarantees deterministic execution. Analysis of the implementation reveals that it uses the implementation of the QM method from the SymPy package (Meurer et al., 2017). To the best of our knowledge, there exists no formal analysis that the execution of SymPy is guaranteed to be deterministic. [^11] The authors report analyzing the code of SymPy and testing its results, aiming at ensuring that the execution is deterministic (McTavish et al., 2025d). However, to the best of our knowledge there exist no formal guarantees that the execution of the set-covering step will be deterministic. As noted earlier, formally ensuring deterministic execution of sequential programs is a challenge in itself Mudduluru et al. (2021); Miao et al. (2025).
The previews arguments and the proof in (McTavish et al., 2025c, a, Theorem 3.4) justify the following result.
Proposition 2
Regarding the use of the QM method for deciding predictive equivalence:
@@
    The QM method can yield incorrect results (with respect to predictive equivalence) if the input to the set-covering step (i.e. the sets of prime implicants and minterms) is not sorted and the implementation of the QM method is not guaranteed to be deterministic.
@@
    The QM method will yield correct results (with respect to predictive equivalence) if the input to the set-covering step is sorted, and the execution of the set-covering step is guaranteed to be deterministic.
Finally, as shown in the next section, the use of the QM method for deciding predictive equivalence is unnecessary. In fact, this paper proves that deciding predictive equivalence is polynomial-time solvable.
@@ 4 Polynomial-Time Algorithms
This section shows that the computational overhead of the approach of McTavish et al. (2025c, a) can be avoided. Concretely, this signifies that the computation of $\mathsf{BCF}{{\mathcal{T}}}$ and $\mathcal{T}_{\mathsf{DNF}}$ (and so the execution of the QM method) is unnecessary for the computational problems studied in McTavish et al.’s work (McTavish et al., 2025c, a). To prove this claim, we show that each algorithm proposed in (McTavish et al., 2025c, a) that uses $\mathcal{T}_{\mathsf{DNF}}$ can be replaced by an algorithm with a running time polynomial in the size of the original DT.
@@ 4.1 Completeness & Succinctness
Given the above, a partial assignment $\mathcal{A}$} is a WAXp of DT $\mathcal{T}$} if and only if it is also a WAXp of any DT $\mathcal{T}$$^{\prime}$ that is predictive equivalent to $\mathcal{T}$}. Furthermore, deciding prediction in the presence of missing data is tightly related with deciding whether a partial assignment represents a WAXp, i.e. it is sufficient for the prediction.
@@
As noted in recent work (Izza et al., 2022; Audemard et al., 2022b; McTavish et al., 2025c, a), decision trees can exhibit path literal redundancy, and so the explanation for a DT prediction can be smaller, possibly much smaller, than the path size in a DT. McTavish et al.’s propose to computing such explanations (and so deciding predictions given missing data) using ${\mathcal{T}}_{\mathsf{DNF}}$. However, as proved in Section˜3 this is worst-case exponential in time and space. In contrast, earlier work proposed a polynomial-time algorithm for explaining an instance $(\mathbf{v},c)$ (e.g. (Izza et al., 2020, 2022)). Here, we opt instead for explaining a partial assignment that is known to be a WAXp for some class $c$. Algorithm˜4 summarizes the computation of one AXp, which requires that the starting partial assignment is an WAXp for class $c$.
@@
The algorithm’s invariant is that the working partial assignment $\mathcal{X}$} is a WAXp, from which features are removed while the invariant is preserved. Due to monotonicity of the definition of WAXp, the partial assignment that is returned represents an AXp. Clearly, the algorithm ensures the property of succinctness of explanations, and runs in polynomial time in the size of the DT. Algorithm˜4 mimics the algorithm proposed in earlier work (Izza et al., 2022), but adapted to working with partial assignments instead of a concrete instance.
@@
Finally, this section shows that the properties of completeness and succinctness are tightly related, and in practice are solved with very similar algorithms, all of which are based on finding/deciding (W)AXps.
@@ 5 Experimental Evidence
The work of McTavish et al. (McTavish et al., 2025c, a) offers comprehensive evidence regarding the importance of eliminating predictive equivalent DTs from the Rashomon set. Also, the algorithms proposed in the previous section exhibit negligible running times, even for DTs of fairly large size. (Our results concur with similar results were obtained in earlier work when computing instance-specific explanations (Izza et al., 2020, 2022).) As a result, the focus of our experiments is to demonstrate the practical effect of exercising the worst-case running time of the QM method. Since the proof Theorem˜1 only considers a single instance, the experiments target the complete compilation of $\mathsf{BCF}$ for different values of $r$.
@@ Results.
Finally, we underscore that the experimental results consider a special DT construction aiming at highlighting the performance limitations of using an implementation of the QM method. However, DTs of large size can occur in some high-risk (Rudin, 2019) application domains, where high accuracy is paramount. For example, (Ghiasi et al., 2020) reports a DT with 161 nodes and 20 real-valued features. In those cases, the use of the QM method will also not scale, even if high-performance solutions are exploited (Coudert, 1994) (and these might not ensure deterministic execution). (We note that SymPy is known to be inefficient for formulas with more than 8 boolean variables. [^15] This is discussed for example in https://docs.sympy.org/latest/modules/logic.html.)
@@ 6.2 Predictive Equivalence and Corrected Measures of Importance
It is well-known that SHAP scores approximate a theoretical definition of Shapley values for XAI (Strumbelj and Kononenko, 2010, 2014). In addition, the quality of the approximation provided by SHAP scores depends on a number of factors. However, recent work proved that the proposed definition of Shapley values for XAI (Strumbelj and Kononenko, 2010, 2014; Lundberg and Lee, 2017) can mislead human decision makers (Marques-Silva and Huang, 2024; Huang and Marques-Silva, 2024), by assigning no importance to features that are critical for a given prediction, and by assigning some importance to features that bear no influence in that prediction. Since it is unclear when Shapley values for XAI (and so SHAP scores) can mislead, the use of the existing definition of Shapley values for XAI should be discouraged.
@@
As argued earlier in this section, in practice appears harder to compute the Shapley values for every instance than simply deciding predictive equivalence. Theorem˜10 shows the connections that logic-based XAI reveals. Furthermore, Theorem˜10 addresses one example of several measures of importance that can be considered:
@@ 6.3 Complexity Results Beyond DTs
Regarding predictive equivalence, this paper (see Algorithm˜5) proves that the problem is in P for decision trees. The same algorithm can be adapted for tractable fragments of multivariate decision trees (Carbonnel et al., 2023; Cooper, 2025). While in general the comparison of representations for predictive equivalence can be unsound if representations are non-canonical, this is a correct approach for when classifiers exhibit canonicity. For example, this is well-known to be the case with $\{\mathsf{BCF}_{c}(C)\,|\,c\in{\mathcal{K}}\}$, i.e. the sets of prime implicants for each predicted class for classifier $C$ (Blake, 1937; Brown, 1990; Darwiche and Marquis, 2002). In addition, it is well-known that reduced ordered binary decision diagrams (ROBDDs) (Bryant, 1986; Darwiche and Marquis, 2002) and sentential decision diagrams (SDDs) (Darwiche, 2011) are canonical representations (Bryant, 1986; Darwiche and Marquis, 2002; Van den Broeck and Darwiche, 2015). As a result, predictive equivalence can be decided in polynomial-time for classifiers represented with ROBDDs and SDDs.